OpenStudy (anonymous):

-4x - 10y = 132 -4x + 2y = 12

OpenStudy (anonymous):

@campbell_st :)

OpenStudy (anonymous):

@campbell_st is my new fav. (:

OpenStudy (campbell_st):

eliminate x by multiplying the 2nd equation by -1 and adding -4x -10y = 132 4x - 2y = -12 --------------- -12y = 120 y = -10 substitute into either equation to find x -4x -10(-10) = 132 x = -8

OpenStudy (anonymous):

8x + 4y = -24 6x - 8y = -84